Express the function f(x)=-2x^3+6x^2-9/2x in factored form and show all work done in order to optain the factored form of the function, also find all zeros of f(x) and state its multiplicity.
Answer by ankor@dixie-net.com(22740)

f(x)=-2x^3+6x^2-9/2x in factored form
write it:
-2x^3 + 6x^2 - 4.5x = 0
Factor out -1, changes the signs, makes it easier to factor
-x(2x^2 - 6x + 4.5) = 0
Factor the quadratic
-x(2x - 3)(x - 1.5) = 0
The zeros
x = 0
and
2x = 3
x = +1.5
and
x = +1.5